### v2

@@ 1 Introduction
Many believe that the constructs echo chamber and filter bubble can explain the decline in democratic exchange, as they cause social media users to become polarized (e.g., Bail et al. [2], Minh Pham et al. [5]) and radicalized. Echo chambers – a term first introduced by Sunstein [6] who repeatedly emphasized the dangers of echo chambers – are frequently defined similarly as “environments in which the opinion, political leaning, or belief of users about a topic gets reinforced due to repeated interactions with peers or sources having similar tendencies and attitudes” [7]. In the context of social media, this involves the interchangeably used but differing concept of filter bubbles. Filter bubbles – coined by Pariser [8] – are created by personalized recommendation systems that expose users to content similar to their beliefs. These algorithms rank and moderate content to provide users with a personalized universe of information [8]. Similarly to previous research, we will use the term echo chamber to describe both Sunstein’s construct of an echo chamber and Pariser’s filter bubble. This is due to the similarity in conceptualization and the fact that Sunstein included algorithms in the characterization of echo chambers in later work (see e.g., Sunstein [9]).
@@
    How does a bibliometric synthesis of existing research characterize the evidence, antecedents, and effects of echo chambers in social networks?
@@
The systematic review offers an overview of antecedents, properties (like existence), and effects of echo chambers in social networks based on research published until December 31st, 2023. We reviewed 112 peer-reviewed studies. From an initial set of 1.651 studies, we selected these 112 studies based on criteria explained in Section 3.5. We identify a taxonomy of conceptualizations and operationalizations of echo chambers and determine antecedents and effects that organize existing work and support future studies. This analysis aims to highlight certain assumptions and choices in the research design of echo chambers. We find that studies differ in (1) outcomes, (2) focus, (3) construct conceptualization and (4) operationalization, and (5) granularity.
@@
Furthermore, we argue that a broader perspective on antecedents, properties, and effects of echo chambers as a concept is essential to understanding the sociotechnical structure and mechanisms of social media. We understand social media and recommendation technologies as sociotechnical systems embedded in a complex, dynamic political and social system that may have multiple feedback loops. A multi-perspective and multi-granular (individual level, structural level) approach to the topic is crucial to promote the understanding of these structures and their reproduction by such systems. This distinguishes this review from existing reviews on echo chambers, echo chamber detection, filter bubbles, and social media.
@@ 3 Method
We follow the PRISMA 2020 [23] guideline to present a transparent account of how we conducted our systematic review. A systematic review following the PRISMA 2020 guideline involves the following elements [24, 25]: (1) a definition of research questions (see below), a strategy regarding the search for relevant literature to answer these questions (Section 3.1), explicit criteria for inclusion and exclusion of research literature (Section 3.4.3), synthesis of evidence that can be derived from the literature (Appendix, Section A), and a summary of the results presented in a structured manner (Section 4).
@@ 3.2 Databases
Social Media* OR Social Network* OR Social Platform* OR Youtube* OR Twitter* OR X* OR Facebook* OR Instagram* OR Tik Tok* OR Reddit* OR YouTube* OR WeChat* OR Weibo* OR QQ* OR Kuaishou* OR Douyin* OR CSSpchat* OR Pinterest* OR WhatsApp* OR Telegram* OR Signal*
@@ 3.3 Search Results
Database searches yielded a total of 1,651 results. Each source’s bibliographic information and abstract were downloaded into a single Citavi library [^2] See https://www.citavi.com/en. Citavi was used to remove duplicates during the inclusion and exclusion process and to code papers in the screening phase. 514 duplicates were removed. Furthermore, only peer-reviewed work was included as justified in Section 3.4. This selection resulted in 996 unique publications that were then screened for eligibility.
@@ 3.4.1 Peer-reviewed Work and Publication Bias
The likelihood of publication bias is one of the possible drawbacks of systematic reviews [27]. Publication bias refers to the tendency of researchers, reviewers, and editors to submit or accept articles for publication based on the direction or intensity of the study findings.
@@ 3.4.2 Other Potential Biases in Systematic Reviews
We want to briefly discuss potential biases that can occur in systematic reviews and how they apply to this one. Location bias refers to the publication of research findings in journals with varying accessibility and levels of indexing in standard databases. For this review, we excluded 42 studies that were inaccessible to us.
First, we address accessibility bias, acknowledging that eleven studies were excluded from this systematic review due to the authors’ limited research access. While our examination of citations suggests that these studies were not particularly influential, we recognize that their exclusion could still introduce bias and potentially overlook valuable perspectives.
Language bias occurs when research findings are published in a specific language based on the type and direction of the results. In this review, we only used English keywords and included studies written in English for transparency and reproducibility reasons. This means that studies relating to certain countries might be excluded from our corpus. This must be taken into account, especially with regard to results relating to certain countries and results relating to research in the global south.
Outcome reporting bias involves selectively reporting some outcomes while omitting others based on the kind and direction of the findings [34]. Due to the large amount of studies, we cannot discuss each in detail. By compiling the outcomes of studies in a table in the Appendix (see A) we aim to give an overview of the studies’ diverse methodologies, results, and conceptualizations.
@@ 3.5 Screening of Search Results
The identification of studies and the screening process in the PRISMA 2020 format can be seen in Figure 1. During the screening process, the eligibility criteria were applied in a hierarchical cascade method, with each source being verified first against C1, then against C2, and then against C3. A total of 1051 studies were screened based on title, excluding 650 studies that did not meet the three criteria. Most papers were excluded based on criterion C1. During the first screening round, we decided if an article was eligible based on its title. This step was conducted through one coder (the first author) who flagged articles to be discussed with a second coder (the second author) to discuss whether an exclusion is warranted. In this phase, only clearly out-of-scope articles, like articles on physical echo chambers, were excluded. In the subsequent screening round, articles were excluded based on their abstract. To calculate Inter-coder reliability (ICR), 20 % of articles were coded by a third coder (the third author) who was unfamiliar with the previous discussions [35, 36]. The agreement between the coding conducted by the first author and the third author – student researcher – was moderate, and differences were then discussed. After the adjustment, the value for Cohens’ $\kappa$ was $0.62$ [37], meaning a substantial agreement.
@@ 3.6 Analysis and Synthesis
The final full-text coding followed an inductive approach to content analysis, as research on echo chambers is dispersed and inconclusive [38]. Codes were established without a preconceived theoretical framework and then grouped into categories concerning conceptualizations and operationalizations of echo chambers, as well as findings concerning antecedents, properties, and effects [39]. An example of a code that was inductively formed is “Focus and Granularity” (see LABEL:tab:coding_subcoding) for all codes and subcodes). During the coding process, it became apparent that there are significant differences between studies in regards to these two aspects. Some studies look at specific groups, others focus on individuals and their behaviourial patterns. Some studies are focused on one platform, others focus on multiple platforms. We wanted to highlight these differences since platform design can influence the behavior of users [40]. Other codes were also formed based on the intention to highlight similarities and differences (e.g. type of data used). A short version of the codebook is presented in Table LABEL:tab:coding_subcoding.
@@
Table 2: Table of the codes and subcodes from the detailed analysis. The more detailed codebook with descriptions is shown in Appendix A.
@@ 4 Results
In this section, we start by giving a brief corpus overview and define conceptualizations and operationalizations that are used across result sections that are ordered by research questions. That means that we follow this brief overview with an analysis of the evidence on the existence of echo chambers, their antecedents, and their effects (RQ1). This is followed by a synthesis of how echo chambers are measured, meaning their conceptualization and operationalization (RQ2). Finally, we analyze the relationships of echo chamber research dimensions to outcomes of research (RQ3). Such analysis includes the relationships between conceptualizations and outcomes, operationalizations and outcomes as well as geographical and contextual factors.
The 129 studies included in the systematic review were all published between 2014 and 2024, as Figure 4 illustrates. The year of publication was not restricted in the search query for this review. There are two possible reasons for these results. First, echo chamber as a term and a topic is relatively new, although some theoretical work was published before 2014. Second, before 2012, there was little research on social media and the impact of social networks on society [43]. The distribution of the identified studies across different publication sources demonstrates the interdisciplinarity of this topic but also represents the need to systemize operationalizations (see Table 3).
In this review, conceptualizations refer to the theoretical frameworks that define and explain the phenomenon. These include homophily—the tendency of individuals to associate with like-minded others, content exposure, which examines the diversity or narrowness of information individuals encounter, and user behavior versus group behavior, which distinguishes between individual-level dynamics and collective patterns within online communities.

### v3

@@ 1 Introduction
Many believe that the constructs echo chamber and filter bubble can explain the decline in democratic exchange, as they cause social media users to become polarized (e.g., Bail et al. [2], Minh Pham et al. [5]) and radicalized (e.g., Torregrosa et al. [6]). Echo chambers – a term first introduced by Sunstein [7] who repeatedly emphasized the dangers of echo chambers – are frequently defined similarly as “environments in which the opinion, political leaning, or belief of users about a topic gets reinforced due to repeated interactions with peers or sources having similar tendencies and attitudes” [8]. In the context of social media, this involves the interchangeably used but differing concept of filter bubbles. Filter bubbles – coined by Pariser [9] – are created by personalized recommendation systems that expose users to content similar to their beliefs. These algorithms rank and moderate content to provide users with a personalized universe of information [9]. Similarly to previous research, we will use the term echo chamber to describe both Sunstein’s construct of an echo chamber and Pariser’s filter bubble. This is due to the similarity in conceptualization and the fact that Sunstein included algorithms in the characterization of echo chambers in later work (see e.g., Sunstein [10]).
@@
    How does academic literature characterize echo chambers in social networks, including their antecedents and effects?
@@
The systematic review offers an overview of antecedents, properties (like existence), and effects of echo chambers in social networks based on research published until December 31st, 2023. We reviewed 129 peer-reviewed studies. From an initial set of 1,706 studies, we selected these 129 studies based on criteria explained in Section 3.5. We identify a taxonomy of conceptualizations and operationalizations of echo chambers and determine antecedents and effects that organize existing work and support future studies. This analysis aims to highlight certain assumptions and choices in the research design of echo chambers. We find that studies differ in (1) outcomes, (2) focus, (3) construct conceptualization and (4) operationalization, and (5) granularity.
@@
Furthermore, we argue that a broader perspective on antecedents, properties, and effects of echo chambers as a concept is essential to understanding the sociotechnical structure and mechanisms of social media. We understand social media and recommendation technologies as sociotechnical systems embedded in a complex, dynamic political and social system that may have multiple feedback loops. A multi-perspective and multi-granular (individual level, structural level) approach to the topic is crucial to promote the understanding of these structures and their reproduction by such systems. This distinguishes this review from existing reviews on echo chambers [15, 21], echo chamber detection [22], filter bubbles [23], and social media [4].
Our review identifies variations in measurement approaches, and regional, political, cultural, and platform-specific biases as key factors contributing to the lack of consensus. Studies based on homophily and computational social science methods often support the echo chamber hypothesis, while research on content exposure that analyzes broader media environments tends to challenge it.
@@ 3 Method
We follow the PRISMA 2020 [26] guideline to present a transparent account of how we conducted our systematic review. A systematic review following the PRISMA 2020 guideline involves the following elements [27, 28]: (1) a definition of research questions (see below), a strategy regarding the search for relevant literature to answer these questions (Section 3.1), explicit criteria for inclusion and exclusion of research literature (Section 3.4.3), synthesis of evidence that can be derived from the literature (Appendix, Section Appendix), and a summary of the results presented in a structured manner (Section 4).
@@ 3.2 Databases
Social Media* OR Social Network* OR Social Platform* OR Youtube* OR Twitter* OR X* OR Facebook* OR Instagram* OR Tik Tok* OR Reddit* OR YouTube* OR WeChat* OR Weibo* OR QQ* OR Kuaishou* OR Douyin* OR Snapchat* OR Pinterest* OR WhatsApp* OR Telegram* OR Signal*
@@ 3.3 Search Results
Database searches yielded a total of 1,706 results. Each source’s bibliographic information and abstract were downloaded into a single Citavi library [^2] See https://www.citavi.com/en. Citavi was used to remove duplicates during the inclusion and exclusion process and to code papers in the screening phase. 514 duplicates were removed. Furthermore, only peer-reviewed work was included as justified in Section 3.4. This selection resulted in 1,051 unique publications that were then screened for eligibility.
@@ 3.4.1 Peer-reviewed Work and Publication Bias
We want to briefly discuss potential biases that can occur in systematic reviews and how they apply to this one. The likelihood of publication bias is one of the possible drawbacks of systematic reviews [30]. Publication bias refers to the tendency of researchers, reviewers, and editors to submit or accept articles for publication based on the direction or intensity of the study findings.
@@ 3.4.2 Other Potential Biases in Systematic Reviews
Location bias refers to the publication of research findings in journals with varying accessibility to researchers and levels of indexing in standard databases. For this review, accessibility bias occured, since we found 11 studies that were inaccessible to us during the screening phase. While these studies have a low citation count, we recognize that their exclusion could still introduce bias and potentially overlook valuable perspectives.
Language bias occurs when research findings are published in a specific language based on the type and direction of the results. In this review, we only used English keywords and included studies written in English for transparency and reproducibility reasons. This means that studies relating to certain countries might be excluded from our corpus and there could be results relating to non-English speaking countries and the Global South that are not included in our analysis. Outcome reporting bias involves selectively reporting some outcomes while omitting others based on the kind and direction of the findings [37]. Due to the large amount of studies, we cannot discuss each in detail. By compiling the outcomes of studies in a table in the Appendix (see Appendix) we aim to give an overview of the studies’ diverse methodologies, results, and conceptualizations.
@@ 3.5 Screening of Search Results
The identification of studies and the screening process in the PRISMA 2020 format can be seen in Figure 1. During the screening process, the eligibility criteria were applied in a hierarchical cascade method, with each source being verified first against C1, then against C2, and then against C3. A total of 1,051 studies were screened based on title, excluding 744 studies that did not meet the three criteria. Most papers were excluded based on criterion C1. During the first screening round, we decided if an article was eligible based on its title. This step was conducted through one coder (the first author) who flagged articles to be discussed with a second coder (the second author) to discuss whether an exclusion is warranted. In this phase, only clearly out-of-scope articles, like articles on physical echo chambers, were excluded. In the subsequent screening round, articles were excluded based on their abstract. To calculate Inter-coder reliability (ICR), 20 % of articles were coded by a third coder (the third author) who was unfamiliar with the previous discussions [38, 39]. The agreement between the coding conducted by the first author and the third author – student researcher – was moderate, and differences were then discussed. After the adjustment, the value for Cohens’ $\kappa$ was $0.62$ [40], meaning a substantial agreement.
@@ 3.6 Analysis and Synthesis
The final full-text coding followed an inductive approach to content analysis, as research on echo chambers is dispersed and inconclusive [41]. Codes were established without a preconceived theoretical framework and then grouped into categories concerning conceptualizations and operationalizations of echo chambers, as well as findings concerning antecedents, properties, and effects [42]. An example of a code that was inductively formed is “Focus and Granularity” (see 2) for all codes and subcodes). During the coding process, it became apparent that there are significant differences between studies in regards to these two aspects. Some studies look at specific groups, others focus on individuals and their behaviourial patterns. Some studies are focused on one platform, others focus on multiple platforms. We wanted to highlight these differences since platform design can influence the behavior of users [43]. Other codes were also formed based on the intention to highlight similarities and differences (e.g. type of data used). A short version of the codebook is presented in Table 2.
@@
Table 2: Table of the codes and subcodes from the detailed analysis. The more detailed codebook with descriptions is shown in Appendix Appendix.
@@ 4 Results
In this section, we start by giving a brief corpus overview and then present the results of our analysis in sections corresponding to the research questions.
The 129 studies included in the systematic review were all published between 2014 and 2024, as Figure 4(a) illustrates. The year of publication was not restricted in the search query for this review. The distribution of the identified studies across different publication sources demonstrates the interdisciplinarity of this topic but also represents the need to systemize operationalizations (see Table 5 in the Appendix).
In this review, conceptualizations refer to the theoretical frameworks that define and explain the phenomenon. These include homophily, which describes the tendency of individuals to associate with like-minded others, content exposure, which examines the diversity or narrowness of information individuals encounter, and user behavior versus group behavior, which distinguishes between individual-level dynamics and collective patterns within online communities.